OpenStudy (lukecrayonz):

Of all numbers whose difference is 18, find the two that have the minimum product

OpenStudy (lukecrayonz):

0 and -18?

OpenStudy (lukecrayonz):

Nope, since you can get negative #'s.. hmmm

OpenStudy (lukecrayonz):

-9 and 9?

OpenStudy (campbell_st):

well let the numbers be x and x - 18 so there product is \[P(x) = x(x -18)....or.....P(x) = x^2 - 18x\] now find the 1st derivative... and then let it equation 0 and solve for x. this will be the minimum value....

OpenStudy (anonymous):

this isnt the exact question but maybe it will give you helpful steps http://answers.yahoo.com/question/index?qid=20080314095143AAmHFLp

OpenStudy (lukecrayonz):

2x-18=0 x=9,-9